OpenStudy (priyar):

if 1+x^4+x^5=sigma(i=0 to 5)a sub i(1+x)^i, for all x in R,then asub2=?

OpenStudy (priyar):

how to go about this problem?

OpenStudy (freckles):

\[1+x^4+x^5\\ =a_0+a_1(1+x)+a_2(1+x)^2+a_3(1+x)^3+a_4(1+x)^4+a_5(1+x)^5\] long way is to expand that ugly thing and combine like terms and then compare coefficients on both sides... let me think about a short way...

OpenStudy (priyar):

ya i thought that would take a lot of time..and didn't proceed

Parth (parthkohli):

Was there a time limit? Was this on a test?

OpenStudy (priyar):

yes

Parth (parthkohli):

Here is one method. Let \(1 + x=t\).\[1 + (t-1)^4 + (t-1)^5 = \sum_{i=0}^5a_it^i\]Easier to expand the left side.

OpenStudy (freckles):

i think you can get the answer by differentiating 3 times and then setting x to -1

OpenStudy (anonymous):

This is high school stuff? o-o

OpenStudy (anonymous):

what grade?

OpenStudy (freckles):

oh wow yeah I got what parth would get

OpenStudy (priyar):

@ParthKohli even now its a bit too long...but i'll try this way

OpenStudy (freckles):

after all if both sides are the same function then their derivatives should be equal and then their double derivatives should be equal and then their triple derivatives should be equal so it makes sense

OpenStudy (freckles):

you will find parth thing easy to expand if you use pascal's triangle

OpenStudy (priyar):

i used binomial theorem

OpenStudy (priyar):

i m getting:\[1+t^5-4t^4+6t^3-4t^2+t\]

OpenStudy (priyar):

on the left side

OpenStudy (freckles):

that looks right

OpenStudy (freckles):

and just compare both sides to find the a_2 value

Parth (parthkohli):

ooo, then we didn't even need to expand the entire thing.

OpenStudy (priyar):

what i did'nt get u?

OpenStudy (freckles):

there is actually a formula for finding binomial coefficients is what he is referring to

OpenStudy (priyar):

anyway it was a nice idea @ParthKohli

OpenStudy (priyar):

i got the answer:-4

OpenStudy (freckles):

my way for fun: \[1+x^4+x^5 \\ = \\ a_0+a_1(1+x)+a_2(1+x)^2+a_3(1+x)^3+a_4(1+x)^4+a_5(1+x)^5 \\ \text{ differentiate both sides 3 \times } \\ \text{ should give } \\ 12x^2+20x^3=2a_2+6a_3(1+x)+12a_4(1+x)^2+20a_5(1+x)^3 \\ \text{ then replace } x=-1 \\ \\ 12(-1)^2+20(-1)=2a_2 \\ \text{ solve for } a_2 \]

Parth (parthkohli):

nice :)

OpenStudy (priyar):

oh nice idea @freckles!

OpenStudy (freckles):

you nice job too I didn't even think about doing the whole t=x+1 thing and using binomial

OpenStudy (priyar):

thank u so much

OpenStudy (priyar):

i have 1 more q..

OpenStudy (priyar):

can any of u help?

OpenStudy (freckles):

what's the question

OpenStudy (priyar):

f(n)=\[[1/3 + 3n/100]n where [n] denotes the gratest integer less than or equal \to n.then \sum_{n=1}^{56} f(n) is ??\]

OpenStudy (priyar):

can u view the full q?

Parth (parthkohli):

are you preparing for jee

Parth (parthkohli):

\[f(n) = n \lfloor \frac{1}3 + \frac{3n}{100}\rfloor \]This is the function, right?

OpenStudy (priyar):

yes

OpenStudy (priyar):

can u read the full Q?

Parth (parthkohli):

Break the bracket part of it (floor/greatest integer function) into pieces: if \(5/3 > 3n/100\ge 2/3 \) then it is 1. and so on.

OpenStudy (priyar):

ya i kind of did it like that but it took some time.. and i had to check like this for each term..

OpenStudy (priyar):

is it the proper way?

Parth (parthkohli):

yes, this is the proper way. just find the intervals where it is zero, one, two and so on.

OpenStudy (priyar):

like that for all the 56 terms?

Parth (parthkohli):

oh, you can't afford to plug in each and every number. just find the \(ranges\).

OpenStudy (priyar):

well..what i did was: i found the ranges .. like..till n=24 the value inside the braket is going to remain<1 right?

OpenStudy (priyar):

then till n=57 the value inside the bracket is gonna be <2..

OpenStudy (priyar):

then i did n(n+1)/2 till n=24 to sum them up

OpenStudy (priyar):

that is 300

OpenStudy (priyar):

then i added from 25,26,......,56 using sum of an A.P... i got 800

OpenStudy (priyar):

so on the whole 1100..but thats not the answer

OpenStudy (priyar):

where did i go wrong?

OpenStudy (priyar):

@freckles ? @ParthKohli ?

OpenStudy (freckles):

\[n \le \lfloor x \rfloor <n+1 \\ n \text{ is an integer } \\ \text{ you only need 3 intervals \to look at } \\ 0 \le \frac{100+9n}{300} <1 \text{ then } \lfloor \frac{100+9n}{300} \rfloor=0 \\ 1 \le \frac{100+9n}{300}<2 \text{ then } \lfloor \frac{100+9n}{300} \rfloor=1 \\ 2 \le \frac{100+9n}{300} <3 \text{ then } \lfloor \frac{100+9n}{300} \rfloor=2\] solve each for n I'm not sure have you done this?

OpenStudy (freckles):

oops i didn't finish my first line \[n \le \lfloor x \rfloor \le n+1 \text{ then } \lfloor x \rfloor=n\]

OpenStudy (priyar):

i got just 2 intervals 0 to 1 and 1 to 2 considering n can take max 56 only...

OpenStudy (freckles):

example first inequality \[0 \le \frac{100+9n}{300} <1 \\ 0 \le 100+9n<300 \\ -100 \le 9n <200 \\ \frac{-100}{9} \le n <\frac{200}{9} \\ \\ \text{ intersect with } 1 \le n \le 56 \\ \text{ we have } 1 \le n \le 22 \\ \text{ so } n \in [1,22] \text{ then } \lfloor \frac{100+9n}{300} \rfloor=0\]

OpenStudy (freckles):

and so on...

OpenStudy (priyar):

oh i stoped at 24...but i must stop at 22..thanks

OpenStudy (freckles):

right you only want to include integer values of n between -100/9 and 200/9 that intersect with integer values in [1,56]

OpenStudy (freckles):

but then you have two more intervals to consider (I say two more as a hint but this should become clear after you do the solving of the inequalities above)

OpenStudy (freckles):

I'm going to go asleep now it is 1:43 am for me

OpenStudy (freckles):

unless you have a question about this question

OpenStudy (priyar):

ok sorry to have disturbed u

OpenStudy (priyar):

thank u so much

OpenStudy (freckles):

no please don't say that you didn't I like math it is just late and I'm old old people get sleepy and they have to work

OpenStudy (freckles):

i liked your questions they were fun (I don't say that about all the questions on here :p)

OpenStudy (priyar):

thank u

OpenStudy (freckles):

so are you good on this problem?

OpenStudy (priyar):

well....i found the intervals..by ur method...[1,22],[22,55] and 56

OpenStudy (priyar):

we can leave [1,22] coz inside the bracket we will have zero

OpenStudy (priyar):

so i added the A.P 23,24.....,55

OpenStudy (freckles):

[1,22] [23,55] {56}

OpenStudy (freckles):

just corrected that one 22 to 23

OpenStudy (priyar):

then atlast i added this to 2*56

OpenStudy (priyar):

is that correct?

OpenStudy (freckles):

you have 3 terms the last term is 2*56

OpenStudy (priyar):

the sum of A.P=1343

OpenStudy (freckles):

I do not get that number

OpenStudy (priyar):

1343+(2*54) is this the answer?

OpenStudy (freckles):

I don't get that either

OpenStudy (freckles):

didn't you say [1,22] we have that floor part is 0 and [23,55] we have that floor part is 1 and then for {56} we have that floor part is 2 correct?

OpenStudy (freckles):

\[\sum_{i=1}^{22} i \cdot 0 + \sum_{i=23}^{55} i \cdot 1 +56 \cdot 2 \]

OpenStudy (freckles):

that first term is just 0

OpenStudy (priyar):

ya thats what i did

OpenStudy (priyar):

final answer:1455 (this is what i got)but answer:1399

OpenStudy (freckles):

1399 is correct

OpenStudy (freckles):

I think you aren't computing middle term right

OpenStudy (priyar):

i did compute..maybe i made a mistake in adding from 23 to 55 using ap..

OpenStudy (freckles):

\[\sum_{i=23}^{55} i \\ \sum_{i-22=1}^{i=55} i \\ \\ \text{ \let } u=i-22 \\ \\ i=55 \text{ then } u=55-22=33 \\ \text{ so we have } \\ \sum_{u=1}^{33} (u+22) =\frac{33(33+1)}{2}+22(33)\]

OpenStudy (priyar):

i got it!!i had made a silly mistake in calculation

OpenStudy (priyar):

Thanks a lot!!

OpenStudy (freckles):

np you guys have fun:)